OpenStudy (anonymous):

Does anyone have an easy way to explain finding the center of an ellipses, and finding the standard equation? google isn't helping...

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=Qm6nCd1R72E

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=H9JKa6pCdfY

OpenStudy (anonymous):

the guy looks like he could snap someone in half with no effort.

OpenStudy (anonymous):

\[Equation = \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{b^2} = 1\] Where the center is at \[(h, k)\]

OpenStudy (anonymous):

I should note that is for if the major axix runs horizontally meaning